# NeuroPhysNet: A FitzHugh-Nagumo-Based Physics-Informed Neural Network Framework for Electroencephalograph (EEG) Analysis and Motor Imagery Classification

### v1

@@ Abstract
Electroencephalography (EEG) is extensively employed in medical diagnostics and brain-computer interface (BCI) applications due to its non-invasive nature and high temporal resolution. However, EEG analysis faces significant challenges, including noise, nonstationarity, and inter-subject variability, which hinder its clinical utility. Traditional neural networks often lack integration with biophysical knowledge, limiting their interpretability, robustness, and potential for medical translation. To address these limitations, this study introduces NeuroPhysNet, a novel Physics-Informed Neural Network (PINN) framework tailored for EEG signal analysis and motor imagery classification in medical contexts. NeuroPhysNet incorporates the FitzHugh-Nagumo model, embedding neurodynamical principles to constrain predictions and enhance model robustness. Evaluated on the BCIC-IV-2a dataset, the framework achieved superior accuracy and generalization compared to conventional methods, especially in data-limited and cross-subject scenarios, which are common in clinical settings. By effectively integrating biophysical insights with data-driven techniques, NeuroPhysNet not only advances BCI applications but also holds significant promise for enhancing the precision and reliability of clinical diagnostics, such as motor disorder assessments and neurorehabilitation planning.
Keywords:
Physics-informed neural networks EEG analysis Motor imagery classification.
@@ 1 Introduction
Electroencephalography (EEG) is a widely used non-invasive technique for measuring the electrical activity of the brain, providing critical insights into various brain functions and disorders. Due to its high temporal resolution and ease of application, EEG has found widespread use in clinical and research settings, such as epilepsy diagnosis, sleep studies, and cognitive monitoring [27]. One particularly significant application of EEG is in Brain-Computer Interfaces (BCIs), where EEG signals are used to establish direct communication pathways between the brain and external devices. Among BCI paradigms, motor imagery (MI) stands out as a prominent example. MI involves the mental rehearsal of movement without actual physical execution, and it plays a crucial role in developing assistive technologies for individuals with motor impairments [5].
@@
Despite the potential of EEG in such applications, analyzing EEG signals remains challenging due to their inherent complexity. EEG signals are characterized by low spatial resolution, high noise levels, and nonstationarity, making it difficult to extract meaningful features [29]. Traditional neural networks, including multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), have been extensively applied to process EEG signals. These models rely on learning complex patterns directly from data, often using handcrafted features or transformations such as spectral power analysis or spatial covariance matrices [22]. While these approaches have shown promise in certain scenarios, they suffer from several limitations. First, traditional methods typically lack the ability to incorporate prior biophysical knowledge about brain dynamics, limiting their interpretability and robustness across varying datasets [3]. Second, they are prone to overfitting in small datasets, which is a common scenario in medical applications [7]. Third, these models often struggle to generalize across subjects due to individual variability in brain dynamics [33].
Physics-Informed Neural Networks offer a promising alternative to address these limitations. The core advantage of PINNs lies in their ability to embed physical laws or domain-specific constraints directly into the learning process, thereby bridging the gap between data-driven methods and mechanistic understanding [8]. By explicitly incorporating physical rules or mathematical formulations within the model, PINNs seamlessly integrate traditional data modeling with physical modeling, enabling neural networks to not only learn complex patterns from data but also adhere to the underlying physical principles, enhancing the scientific validity of predictions.
In the context of EEG signal analysis, PINNs demonstrate unique advantages. The generation of EEG data is fundamentally governed by neurodynamical mechanisms, such as the variations in neuronal membrane potentials and the propagation of signals across nodes [32]. Traditional data-driven approaches often overlook these inherent physical constraints, while PINNs leverage neurodynamical models to capture these dynamics. These equations, known for their simplicity yet powerful functionality, describe the changes in neuronal membrane potential and recovery dynamics, providing a biophysically consistent representation of neural activity.
Moreover, the ability of PINNs to incorporate biophysical modeling enhances both the interpretability and generalizability of the model. Conventional machine learning methods commonly suffer from issues such as overfitting on small datasets and limited transferability across different datasets [6]. By embedding physical constraints, PINNs introduce a strong inductive bias, reducing dependency on large-scale labeled data and improving robustness against noise and shifts in data distribution [8]. Importantly, by linking the outputs of the neural network directly to validated physical models, PINNs enable researchers to interpret the decision-making process of the model more effectively, a feature that is particularly crucial in medical applications where reliability is paramount.
Building on this foundation, this study introduces a novel PINN-based framework specifically designed for the analysis of EEG signals, with a particular focus on motor imagery classification. This framework integrates the FitzHugh-Nagumo model [11] into the learning process, leveraging its neurodynamical principles to constrain the predictions. By embedding this biophysical model, the framework ensures that the outputs adhere to well-established physiological laws, enhancing the reliability and interpretability of the results. In addition to incorporating the FitzHugh-Nagumo model, the framework includes a dedicated feature extraction module that processes the outputs of the PINN. This module extracts robust temporal features from the biophysically informed variables, such as membrane potentials and recovery dynamics, and transforms them into compact and discriminative representations suitable for downstream classification tasks. This step bridges the gap between physical modeling and modern classification techniques, ensuring that the extracted features retain both local temporal information and global biophysical relevance. By seamlessly combining biophysical modeling and data-driven neural network techniques, the proposed framework effectively addresses the limitations of traditional approaches. It mitigates the challenges of poor generalizability and limited interpretability, often observed in standard neural networks applied to EEG analysis. Furthermore, this integration allows the model to capture essential neurodynamical patterns while remaining flexible enough to adapt to the complex, nonlinear nature of EEG signals. As a result, the framework offers a more interpretable and generalizable solution for EEG-based applications
The primary contributions of this work are as follows:
@@
  • Application to Motor Imagery Classification: We demonstrate the effectiveness of the proposed framework in motor imagery classification tasks, achieving improved performance and generalization compared to traditional neural network approaches. The results underscore the potential of combining biophysical modeling with modern neural network architectures in medical applications.
@@ 2 Related Work
2.1 Deep Learning in BCI-MI Analysis
Recent advances in deep learning have substantially enhanced the performance of BCIs for MI analysis, particularly through the use of CNNs and Recurrent Neural Networks (RNNs). These techniques have demonstrated notable improvements in classification accuracy for MI tasks, as shown in several key studies.
@@ 2.1 Deep Learning in BCI-MI Analysis
Schirrmeister et al. demonstrated the potential of CNNs in EEG decoding and visualization, showcasing the ability of deep learning models to extract meaningful features from EEG signals for improved classification[25]. Similarly, Lawhern et al. introduced EEGNet, a compact CNN architecture specifically designed for efficient EEG-based BCI applications, further advancing the state of the art in MI classification[17]. Huang et al. also investigated CNN-based deep learning models for MI classification, highlighting their superior performance in classification accuracy compared to traditional methods[9]. The low SNR in MI EEG signals poses a challenge in decoding movement intentions, but this can be addressed using multi-branch CNN modules that learn spectral-temporal domain features, as suggested in [13]. Futhermore, Ju et al. proposed Tensor-CSPNet[16] and Graph-CSPNet[14], a novel geometric deep learning framework for motor imagery classification, achieving improved feature extraction, robustness, and interpretability.
Despite their success, deep learning approaches for MI analysis face several inherent limitations. First, they are heavily data-dependent, requiring large labeled datasets for training, which is often not feasible in medical applications with limited data. This leads to issues of overfitting, particularly when working with small datasets. Furthermore, deep learning models are often criticized for their "black box" nature, which hinders their interpretability and transparency—crucial factors for medical applications. Additionally, due to inter-subject variability and differences in brain dynamics, deep learning models often struggle with generalization, limiting their ability to perform reliably across diverse datasets. Lastly, complex neural network architectures tend to result in long training and inference times, making them less suitable for real-time applications in BCIs.
2.2 Physics-Informed Neural Networks in BCI-MI Analysis
To address the aforementioned limitations of traditional deep learning models, PINNs have emerged as a promising alternative. PINNs combine data-driven deep learning methods with physical constraints derived from underlying scientific principles, offering enhanced interpretability, generalization, and robustness.
Raissi et al. first introduced the concept of PINNs as a method for solving forward and inverse problems involving nonlinear partial differential equations (PDEs). Their work demonstrated how incorporating physical laws directly into the learning framework can improve the accuracy and generalization of deep learning models. This approach has since been extended to various domains, including the study of dynamic systems in neurodynamics and brain modeling, which are directly relevant to EEG signal analysis in BCI applications. In particular, PINNs offer a powerful tool for modeling the brain’s biophysical dynamics during motor imagery tasks[1].
Lu et al. explored the use of Physics-Informed DeepONets, an extension of PINNs, for solving parametric PDEs. Their work highlights the ability of PINNs to generalize across different parametric settings, a feature that could be applied to EEG signal analysis, where individual differences in brain activity need to be accounted for. The inclusion of such models could improve the performance of BCIs by better capturing the underlying neurodynamical processes during MI tasks[31]. Zhang et al. proposed a framework to quantify uncertainty in PINNs, which is particularly beneficial for handling noisy EEG data and improving the robustness of models, as uncertainty quantification is crucial in BCI applications where data quality is often variable[34].
Further advancements have been made in improving the convergence and efficiency of PINNs. Jagtap et al. introduced adaptive activation functions that accelerate the training process, reducing the computational burden and improving the model’s performance. This innovation is particularly relevant in the context of real-time BCI applications, where fast and accurate predictions are necessary for immediate feedback to users[12]. Chen et al. applied PINNs to heat transfer problems, demonstrating the potential of such methods in improving model accuracy by incorporating physical constraints. Although their focus was not on EEG signals, the general principle of embedding physical models into the neural network has significant implications for brain signal processing[4].
In the context of long-term integration of parametric ordinary differential equations (ODEs), Wang et al. explored how PINNs could be applied to such problems, providing insights into modeling dynamic systems over extended periods. This capability is essential for BCI-MI systems, where the brain’s activity during motor imagery may evolve over time and require long-term temporal modeling[30]. The application of PINNs in fluid dynamics, as demonstrated by Mao, showcases how complex physical phenomena can be captured through neural networks. This approach can be extended to model the complex dynamics of brain activity, where non-linear interactions between different brain regions play a key role in motor imagery tasks[24].
Our proposed method, which integrates the FitzHugh-Nagumo model within a PINN framework, addresses several of the limitations associated with both deep learning and traditional PINN approaches. By embedding biophysical knowledge directly into the learning process, our model offers enhanced interpretability, a crucial factor for medical applications. The incorporation of physical constraints also reduces the model’s dependence on large labeled datasets, improving generalization across individuals and datasets. Furthermore, the robustness of the approach is enhanced, as the physical constraints allow the model to better handle noise and individual variations in EEG signals. The integration of data-driven techniques with biophysical models in our framework ensures that the model’s predictions are aligned with established physiological laws, providing a more scientifically grounded approach to EEG signal analysis in BCI applications.
@@ 3 Methodology
In this study, we develop a PINN based on the FitzHugh-Nagumo (FHN) model to analyze and interpret EEG data for BCI applications. The proposed architecture seamlessly integrates domain-specific physiological laws with advanced deep learning techniques, thereby enhancing predictive performance and ensuring adherence to biophysical consistency. This section outlines the mathematical foundations, architectural design, and implementation details of the FHN-informed PINN framework.
@@ 3.1 FitzHugh-Nagumo Model
The FHN model stands as a seminal framework in the study of neuronal dynamics, offering a simplified yet robust representation of the essential electrical activities within neurons. Introduced by Richard FitzHugh and Julius Nagumo in the 1960s, the FHN model was conceived to streamline the more intricate Hodgkin-Huxley model while retaining the fundamental characteristics of neuronal action potentials. The Hodgkin-Huxley model details the ionic mechanisms underlying action potentials, providing a precise description of neuronal electrical activity. However, its high dimensionality and complexity pose significant challenges for large-scale network simulations and theoretical analyses. In contrast, the FHN model reduces the number of variables and parameters, presenting a more tractable form to describe the excitation and recovery processes of neurons, thereby facilitating more efficient and feasible analyses and simulations of neural dynamics.
Mathematically, the FHN model is articulated through a pair of coupled nonlinear ordinary differential equations that describe the temporal evolution of two critical variables: the activation variable $u$ and the recovery variable $v$. These equations are given by:
@@
Despite its reduced complexity, the FHN model adeptly captures the cyclical nature of neuronal firing, encompassing the transition from a quiescent state to an excited state and back to rest. This cyclical behavior is emblematic of the generation and propagation of action potentials in real neurons. The FHN model exhibits a rich array of dynamical phenomena, including limit cycle oscillations, bifurcations, and chaotic dynamics, contingent upon the values of the parameters $\epsilon$, $a$, and $b$. Under varying parameter regimes, the model can display distinct behaviors such as stable fixed points (resting states), periodic oscillations (reflecting repetitive firing of action potentials), and chaotic dynamics (indicating complex, aperiodic neuronal activity). These characteristics render the FHN model not only a valuable theoretical tool for understanding neuronal behavior but also a practical model for simulating and analyzing complex neural systems.
In the context of EEG modeling, the FHN model offers several compelling advantages that align with the intricate dynamics observed in brain activity. EEG is a non-invasive technique that records electrical activity generated by large populations of neurons, providing rich spatiotemporal information about brain function. The signals captured by EEG reflect the synchronous firing and oscillatory behavior of neuronal assemblies across different brain regions, encompassing a wide range of frequency bands associated with various cognitive and physiological states.
One of the primary motivations for employing the FHN model in EEG modeling is its ability to encapsulate the core dynamical features of neuronal activity with a minimal set of variables and parameters. The simplicity of the FHN model facilitates the analysis and simulation of large-scale neural networks, which is essential for interpreting EEG data that arise from the collective behavior of thousands to millions of neurons. By reducing the dimensionality of the problem, the FHN model allows for more tractable mathematical and computational treatments, enabling researchers to focus on the essential dynamics without being overwhelmed by the complexity of the underlying biological processes.
Moreover, the FHN model’s capacity to represent both excitatory and inhibitory interactions makes it particularly suited for capturing the balance of neural excitation and inhibition that is crucial for normal brain function. This balance is reflected in EEG signals, where different oscillatory patterns emerge from the interplay between excitatory and inhibitory neural populations. The nonlinear dynamics of the FHN model are adept at reproducing these oscillatory patterns, providing a mechanistic basis for understanding the generation of EEG rhythms such as alpha, beta, and gamma waves.
Another significant advantage of the FHN model in EEG modeling lies in its flexibility and scalability. The model can be easily extended to incorporate additional complexities, such as spatial interactions between neurons or the inclusion of external inputs that mimic sensory stimuli or cognitive tasks. This adaptability is particularly beneficial for applications in BCI, where EEG signals are utilized to decode and interpret specific mental states or intentions, such as motor imagery. The FHN model can be integrated into advanced computational frameworks, such as PINNs, to enhance the accuracy and robustness of EEG signal interpretation by embedding biophysical constraints directly into the learning process.
Furthermore, the FHN model serves as a bridge between theoretical neuroscience and practical applications in medical technology. By providing a clear and interpretable model of neuronal dynamics, the FHN framework facilitates the development of algorithms capable of accurately decoding EEG signals in real-time applications. This is particularly relevant for BCI systems aimed at assisting individuals with motor impairments, where precise and reliable interpretation of EEG data is paramount for effective control of prosthetic devices or communication aids.
In addition to its theoretical and practical utility, the FHN model’s ability to generate simulation data that can be compared with experimental EEG recordings offers a robust mechanism for validating and refining signal processing algorithms. By simulating EEG signals under various conditions and tasks, researchers can systematically evaluate the performance of different modeling approaches and optimize parameters to achieve better alignment with empirical data. This iterative process enhances the reliability and validity of EEG-based models, ensuring that they accurately reflect the underlying neural dynamics.
In conclusion, the FHN model, with its balanced combination of simplicity and dynamical richness, emerges as an indispensable tool for EEG modeling in the realm of BCI applications. Its ability to capture the fundamental aspects of neuronal excitability and inhibition, coupled with its computational efficiency and scalability, provides a solid foundation for developing sophisticated models that can interpret complex brain signals with high accuracy. The integration of the FHN model into frameworks like PINNs not only enhances predictive performance but also ensures the physical consistency of model outputs, thereby bridging the gap between theoretical neuroscience and practical medical applications. As such, the FHN model significantly contributes to advancing our understanding of neural dynamics and driving the development of effective BCI technologies for medical rehabilitation and human-machine interfacing.
@@ 3.2 Design and Implementation of a Physics-Informed Neural Network Based on the FitzHugh-Nagumo Model
In this study, we present the design and implementation of a PINN architecture grounded in the FHN model for the analysis and interpretation of EEG data. The proposed PINN architecture is structured into three components: the “Data Preprocessing Module”, the “Feature Extraction Module”, and the “PINN Model”. This hierarchical design facilitates the transformation of raw EEG signals into a structured format, the extraction of pertinent features, the incorporation of biophysical constraints through the FHN model, and the subsequent classification of specific tasks such as motor imagery. The integration of these modules ensures enhanced predictive performance, robustness, and adherence to the underlying physical principles governing neuronal dynamics.
@@ 3.2.1 Data Preprocessing Module
where W represents the number of window slices, C corresponds to the number of EEG channels, and $\omega$ signifies the window length. The segmentation aims to divide EEG signals into small segments on the time domain, either with overlapping or without overlapping. This division facilitates the detailed analysis of temporal dynamics within each window, allowing the model to capture transient features that may be indicative of specific cognitive or motor states.
@@
where $F$ denotes the number of filter banks applied. Utilizing causal Chebyshev Type II filters, the raw oscillatory EEG signals are decomposed into distinct frequency bands. This decomposition is essential for isolating frequency-specific information corresponding to various neural oscillations, such as alpha, beta, and gamma waves, each associated with different cognitive and physiological functions. By separating the EEG signals into these frequency bands, the model can more effectively capture the nuanced patterns and interactions that occur within specific frequency ranges.
@@ 3.3 Physics-Informed Neural Network Module
The PINN module is the central component of the proposed architecture, responsible for integrating the FHN model’s biophysical constraints with the structured EEG data. This integration ensures the model’s predictions adhere to the underlying physical principles of neuronal dynamics, providing accurate and interpretable results.
@@
The model accepts input $\mathbf{X}^{\prime\prime}$ in the shape $B\times W\times F\times C\times\omega$. To align the input data with the convolutional layers, the tensor is reshaped into $(B\times W)\times F\times C\times\omega$, effectively merging the batch and window dimensions. This reshaped tensor is passed through two 2D convolutional layers, each followed by batch normalization and ReLU activation to stabilize the learning process and introduce non-linearity. These layers capture spatiotemporal patterns in the EEG data, which are crucial for modeling neuronal dynamics. The operations are mathematically represented as:
@@
[…]
@@
Compared to traditional PINNs, which primarily rely on MLPs, the PINN model proposed in this study incorporates a combination of CNNs and Transformer encoders. This design not only offers significant advantages in processing complex EEG data but also demonstrates superior capabilities in solving the key variables $v$ and $w$ of the FHN model.
Traditional PINNs with MLP-based architectures often struggle to effectively capture local spatiotemporal features and long-range dependencies in high-dimensional time-series data. MLPs process input data through fully connected layers without leveraging the spatial and temporal structures of the data. This lack of structural awareness can lead to inefficient feature extraction and reduced accuracy, particularly in estimating the activation variable $v$ and recovery variable $w$.
In this study, CNNs are introduced as the feature extraction module to capture local spatiotemporal features from EEG signals. By using the convolutional kernels, CNN layers scan input signals along the temporal dimension to identify localized patterns and variations. These localized features are critical for accurately modeling the dynamics of $v$ and $w$ in the FHN model. Moreover, the application of Batch Normalization and ReLU activation functions enhances the stability of model training and improves the representation of non-linear features, further benefiting the calculation of $v$ and $w$.
Additionally, Transformer encoders are integrated into the PINN model to capture long-range temporal dependencies. EEG signals often exhibit correlations over extended time spans, which are challenging for traditional convolutional operations to capture effectively. Transformers utilize multi-head self-attention mechanisms to dynamically evaluate relationships across different time points, enabling the model to construct global dependencies. This mechanism provides richer temporal context, ensuring that $v$ and $w$ are calculated with higher physical consistency and dynamic precision.
By combining CNNs and Transformers, the proposed model achieves efficient extraction of both local and global features while ensuring the accurate computation of $v$ and $w$ under the constraints of the FHN equations. This design significantly outperforms traditional MLP-based PINNs by producing results that are more robust and precise when applied to complex EEG data. Furthermore, the inclusion of biophysical constraints ensures that the predicted $v$ and $w$ are not only accurate but also aligned with the underlying biological dynamics, enhancing the interpretability and reliability of the predictions.
@@
[…]
@@
This physics-based constraint ensures that the predictions follow the temporal dynamics governed by the FHN equations, enhancing the interpretability and reliability of the model.
Advantages
This formulation of the physics-based loss:
  • Ensures that the predictions $v$ and $w$ adhere to the FHN model’s dynamics, maintaining physical consistency.
  • Accounts for interactions among nodes through the coupling term, enabling a realistic representation of network dynamics.
  • Uses finite differences to compute time derivatives, reducing computational complexity compared to backpropagation through time.

### v2

@@ Abstract
Electroencephalography (EEG) is extensively employed in medical diagnostics and brain-computer interface (BCI) applications due to its non-invasive nature and high temporal resolution. However, EEG analysis faces significant challenges, including noise, nonstationarity, and inter-subject variability, which hinder its clinical utility.This study introduces NeuroPhysNet, a novel Physics-Informed Neural Network (PINN) framework tailored for EEG signal analysis and motor imagery classification in medical contexts. NeuroPhysNet incorporates the Fitz Hugh-Nagumo model, embedding neurodynamical principles to constrain predictions and enhance model robustness. Evaluated on the BCIC-IV-2a dataset, the framework achieved superior accuracy and generalization compared to conventional methods, especially in data-limited and cross-subject scenarios, which are common in clinical settings. By effectively integrating biophysical insights with data-driven techniques, NeuroPhysNet not only advances BCI applications but also holds significant promise for enhancing the precision and reliability of clinical diagnostics.
@@ 1 Introduction
Electroencephalography (EEG) is a widely used non-invasive technique for measuring the electrical activity of the brain, providing critical insights into various brain functions and disorders. Due to its high temporal resolution and ease of application, EEG has found widespread use in clinical and research settings, such as epilepsy diagnosis, sleep studies, and cognitive monitoring [tatum2018clinical]. One particularly significant application of EEG is in Brain-Computer Interfaces (BCIs), where EEG signals are used to establish direct communication pathways between the brain and external devices. Among BCI paradigms, motor imagery (MI) stands out as a prominent example. MI involves the mental rehearsal of movement without actual physical execution, and it plays a crucial role in developing assistive technologies for individuals with motor impairments [dickstein2007motor].
@@
Despite its potential in various applications, EEG analysis remains challenging due to the signals’ low spatial resolution, high noise levels, and nonstationarity, which complicate the extraction of meaningful features [wang2021review]. Traditional neural networks, such as multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), have been widely used to process EEG signals by learning complex patterns from data, often with handcrafted features or transformations like spectral power analysis or spatial covariance matrices [lu2023lgl]. While effective in some contexts, these approaches face key limitations: they typically lack integration of biophysical priors, reducing interpretability and robustness across datasets [bijsterbosch2020challenges]; they are prone to overfitting on small medical datasets [garcia2023machine]; and they struggle to generalize across subjects due to inter-individual variability in brain dynamics [yu2024learning].
Physics-Informed Neural Networks offer a promising alternative to address these limitations. The core advantage of PINNs lies in their ability to embed physical laws or domain-specific constraints directly into the learning process, thereby bridging the gap between data-driven methods and mechanistic understanding [hao2022physics]. PINNs capture EEG’s neurodynamical basis—membrane potential fluctuations and signal propagation—via concise biophysical equations, addressing the physiological constraints often ignored by purely data-driven methods [yang2020neuromorphic]. Embedding these constraints imparts a strong inductive bias that curbs overfitting on limited datasets, enhances robustness to noise and distributional shifts [dou2023machine, hao2022physics], and affords transparent decision pathways essential for reliable medical applications.
Building on prior PINN advances, we introduce a novel framework for EEG‐based motor imagery classification that embeds the FitzHugh–Nagumo model [izhikevich2006fitzhugh] to enforce neurodynamical constraints and ensure physiologically consistent, interpretable predictions. The primary contributions of this work are as follows:
@@
  • Application to Motor Imagery Classification: We demonstrate the effectiveness of the proposed framework in motor imagery classification tasks, achieving improved performance and generalization compared to traditional neural network approaches.
@@ 2 Related Work
Recent advances in deep learning have substantially enhanced the performance of BCIs for MI analysis, particularly through the use of CNNs and Recurrent Neural Networks. Schirrmeister et al. demonstrated the potential of CNNs in EEG decoding and visualization, showcasing the ability of deep learning models to extract meaningful features from EEG signals for improved classification[2017Deep]. Similarly, Lawhern et al. introduced EEGNet, a compact CNN architecture specifically designed for efficient EEG-based BCI applications[2018EEGNet]. Huang et al. also investigated CNN-based deep learning models for MI classification, highlighting their superior performance in classification accuracy[2022EEG]. The low SNR in MI EEG signals poses a challenge in decoding movement intentions, but this can be addressed using multi-branch CNN modules that learn spectral-temporal domain features, as suggested in [jia2023model]. Futhermore, Ju et al. proposed Tensor-CSPNet[2023Tensor] and Graph-CSPNet[ju2022graph], a novel geometric deep learning framework for motor imagery classification, achieving improved feature extraction, robustness, and interpretability.
Despite their advances in MI analysis, deep learning approaches remains data-hungry and overfits with scarce labels; its black-box nature hampers interpretability; inter-subject and temporal variability limits generalization; and heavy models slow training and inference, hindering real-time BCI.
To address the limitations of traditional deep learning models, PINNs have emerged as a promising alternative. Raissi et al. first introduced PINNs for solving forward and inverse nonlinear PDEs by embedding physical laws into the learning process; this framework was later adapted to neurodynamics and brain modeling for EEG analysis in BCI applications [2019Physics]. Building on this, Lu et al. developed Physics-Informed DeepONets to handle parametric PDEs and individual variability [wang2021learning], while Zhang et al. incorporated uncertainty quantification to bolster robustness against noisy EEG data [zhang2019quantifying]. Finally, Jagtap et al. accelerated convergence with adaptive activation functions—crucial for real-time BCI feedback [jagtap2020adaptive]
Our proposed method, the integration of FitzHugh–Nagumo model into a PINN framework, addresses deep‐learning and standard PINN limitations by embedding biophysical constraints that enhance interpretability, reduce reliance on large labeled datasets, improve generalization across subjects, and bolster robustness to EEG noise and individual variability.
@@ 3 Methodology
In this study, we develop a PINN based on the FitzHugh-Nagumo (FHN) model to analyze and interpret EEG data for BCI applications.
@@ 3.1 FitzHugh-Nagumo Model
The FHN model, introduced by FitzHugh and Nagumo in the 1960s, offers a two-variable reduction of the Hodgkin–Huxley equations that preserves core action-potential features. By abstracting detailed ionic currents into excitation and recovery variables, it overcomes the high dimensionality and computational burden of the full Hodgkin–Huxley framework, enabling tractable, large-scale simulations and theoretical analyses of neuronal dynamics. Mathematically, the FHN model is articulated through a pair of coupled nonlinear ordinary differential equations that describe the temporal evolution of two critical variables: the activation variable $u$ and the recovery variable $v$. These equations are given by:
@@ 3.2 Design and Implementation of a Physics-Informed Neural Network Based on the FitzHugh-Nagumo Model
We propose a three-module PINN architecture—Data Pre-processing, Feature Extraction, and PINN Model—grounded in the FitzHugh–Nagumo equations for EEG analysis.
@@ 3.2.1 Data Preprocessing Module
where W represents the number of window slices, C corresponds to the number of EEG channels, and $\omega$ signifies the window length. The segmentation aims to divide EEG signals into small segments on the time domain, either with overlapping or without overlapping.
@@
where $F$ denotes the number of filter banks applied. Utilizing causal Chebyshev Type II filters, the raw oscillatory EEG signals are decomposed into distinct frequency bands.
@@ 3.3 Physics-Informed Neural Network Module
The PINN module is the central component of the proposed architecture, responsible for integrating the FHN model’s biophysical constraints with the structured EEG data.
@@
The model accepts input $\mathbf{X}^{\prime\prime}$ in the shape $B\times W\times F\times C\times\omega$. To align the input data with the convolutional layers, the tensor is reshaped into $(B\times W)\times F\times C\times\omega$, effectively merging the batch and window dimensions. This reshaped tensor is passed through two 2D convolutional layers, each followed by batch normalization and ReLU activation to stabilize the learning process and introduce non-linearity. The operations are mathematically represented as:
@@
The output tensor $\mathbf{O}$ is reshaped to [batch_size, sliding_windows, 2 $\times$ num_nodes, data_points], where num_nodes represents the number of nodes in the graph, and data_points denotes the feature dimensions for each node. The first num_nodes along the node dimension is extracted as $v$, while the remaining num_nodes is extracted as $w$:
$\mathbf{v}=\mathbf{O}[:,:,:\text{num\_nodes},:]\quad\mathbf{w}=\mathbf{O}[:,:,\text{num\_nodes}:,:]$ (12)
@@
Unlike traditional PINNs based on MLPs, our model employs CNNs and Transformer encoders to exploit EEG’s spatiotemporal structure and accurately estimate the FitzHugh–Nagumo variables $v$ and $w$. Temporal CNNs extract local patterns, stabilized by Batch Normalization and ReLU, while multi-head self-attention captures long-range dependencies to provide global context.
@@
where coupling strength is set to $0.1$, and the coupling matrix $K$, with a shape of $\text{num\_nodes},\text{num\_nodes}]$, is initialized as an all-ones matrix with its diagonal elements subtracted by 1 and then multiplied by the coupling strength.